itzpxyton:

Solve for x: two sevenths(x − 2) = 4x x = two thirteenths x = negative two thirteenths x = negative thirteen halves x = thirteen halves

Shadow:

\[\frac{ 2 }{ 7 }(x - 2) = 4x\]

itzpxyton:

2/7*x = 2/7x

Shadow:

Yes, and it's the distributive property so we're also doing 2/7 times -2, what do you get?

itzpxyton:

yes but i got 0.5 something

Shadow:

\[\frac{ 2 }{ 7 } \times 2 = \frac{ 4 }{ 7 }\] So we're at: \[\frac{ 2 }{ 7 }x - \frac{ 4 }{ 7 } = 4x\]

itzpxyton:

1 attachment
Shadow:

I like that solution a lot better, more pretty. It leaves the 7 under...

itzpxyton:

so would my answer be b

itzpxyton:

x = negative two thirteenths -2/13=x

Shadow:

\[\frac{ 2 }{ 7 } \times 2 = \frac{ 4 }{ 7 }\] So we're at: \[\frac{ 2 }{ 7 }x - \frac{ 4 }{ 7 } = 4x\] \[-\frac{4 }{ 7 } = \frac{ 28 }{ 7 }x - \frac{ 2 }{ 7 }x\] \[-\frac{ 4 }{ 7 } = \frac{ 26 }{ 7 }x\] Multiply both sides by 7 \[-4 = 26x\] \[x = - \frac{ 4 }{ 26 }\] Simplifying \[x = - \frac{ 2}{ 13 }\]

Shadow:

You're right. I was just writing out a full explanation in case anyone else comes across this question.

itzpxyton:

okay

itzpxyton:

Thank you!